Copper has a density of 9 g/cm3 and a bulk modulus of 1.2 x 1011 N/m2. What is the speed of a wave through it?


1
Expert's answer
2022-03-06T15:17:40-0500

The speed of sound

"v_s=\sqrt{B/\rho}\\
=\sqrt{1.2*10^{11}/9000}=3651\:\rm m/s"